### v1

@@ Abstract
Image restoration involves recovering a high-quality clean image from its degraded version, which is a fundamental task in computer vision. Recent progress in image restoration has demonstrated the effectiveness of learning models capable of addressing various degradations simultaneously, i.e., the All-in-One image restoration models. However, these existing methods typically utilize the same parameters facing images with different degradation types, which causes the model to be forced to trade off between degradation types, therefore impair the total performance. To solve this problem, we propose HAIR, a Hypernetworks-based plug-in-and-play method that dynamically generated parameters for the corresponding networks based on the contents of input images. HAIR consists of 2 main components: Classifier (Cl) and Hyper Selecting Net (HSN). To be more specific, the Classifier is a simple image classification network which is used to generate a Global Information Vector (GIV) that contains the degradation information of the input image; And the HSNs can be seen as a simple Fully-connected Neural Network that receive the GIV and output parameters for the corresponding modules. Extensive experiments shows that incorporating HAIR into the architectures can significantly improve the performance of different models on image restoration tasks at a low cost, although HAIR only generate parameters and haven’t change these models’ logical structures at all. With incorporating HAIR into the popular architecture Restormer, our method obtains superior or at least comparable performance to current state-of-the-art methods on a range of image restoration tasks. Code and pre-trained checkpoints are available $\underline{\textbf{Code and pre-trained checkpoints are available here.}}$
@@ 1 Introduction
Image restoration stands as a pivotal challenge within the field of computer vision, dedicated to the reconstruction of high-quality images from their degraded states. The presence of adverse conditions such as noise, haze, or rain can severely diminish the effectiveness of images for a variety of applications, such as autonomous navigation [1, 2], augmented reality [3, 4, 5]. Developing robust image restoration methods is thus an imperative pursuit. The utilization of deep learning into this domain has yielded remarkable progress, as evidenced by a suite of recent methodologies [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Nonetheless, the predominant approach in current research is to employ task-specific models, each tailored to address a particular type of degradation [16, 7, 12, 13, 8]. This tailored approach, while precise, presents a constraint in terms of versatility, as it restricts the applicability of models to scenarios with varied or unknown degradations [17, 18, 19].
To overcome this limitation, the research community has increasingly turned its attention to the development of All-in-One image restoration models. These models are designed to adeptly manage a spectrum of degradations in a single framework. Pioneering efforts in this area [20, 21, 1, 22, 23, 24, 25, 26, 15] have utilized a variety of advanced techniques such as contrastive learning [20], meta-learning [25], visual prompting methods [24, 27, 28, 14]. These approaches have undoubtedly made substantial contributions to the field, propelling it forward with notable advancements.
@@
Figure 1: The difference between our method and previous methods. (a) Previous All-in-One image restoration methods. For images with various degradations, these methods utilize a single model with the same parameters. (b) Our proposed HAIR. Given a certain degraded image, we first use a fixed Extraction module to extract the features, and feed the extracted feature into the Hypernetwork to generate the dynamic parameters for the Fusion module, and finally get the restored image by the generated parameters.
@@
Despite the notable successes of these methods, they share a common drawback: they rely on a single model with fixed parameters to address a variety of degradations. This one-size-fits-all approach can hinder the model’s effectiveness when dealing with multiple degradations simultaneously. For instance, when viewed through the lens of frequency domain analysis, haze is characterized as low-frequency noise, in contrast to rain, which is considered high-frequency interference. An effective dehazing model acts as a low-pass filter, preserving high-frequency details, whereas deraining requires the opposite—enhancing the high-frequency components. Consequently, a model must strike a balance between these conflicting demands of different degradations, which may lead to suboptimal performance.
@@
To mitigate the aforementioned issue, we propose HAIR in this paper. Our fundamental concept is the dynamic generation of parameters based on the degradation information present in the input image. HAIR employs Hypernetworks [29], a trainable neural network that outputs parameters for other neural networks, to take the degradation information from the input image and produce corresponding parameters. Specifically, for a given unknown degraded image, we first utilize a Classifier, similar to those used in image classification networks, to obtain its Global Information Vector. This vector is then passed to the Hyper Selecting Net, which is a Fully Connected Network, to generate the requisite parameters, as illustrated in Fig. 1. Thereby, our method can dynamically adapt to various degradations. With these dynamically-parameterized modules, we ultimately achieve the restored image. In brief, our contributions include:
  • We propose HAIR, a novel Hypernet-based blind All-in-One image restoration method capable of dynamically generating parameters based on the input image’s characteristics. By incorporating HAIR with Restormer [12], our method achieves performance that is superior or at least comparable to that of current state-of-the-art methods across a variety of image restoration tasks.
  • Our proposed HAIR, which comprises a Classifier and Hyper Selecting Nets, functions as a plug-in-and-play module. It can be effortlessly integrated into any existing restoration network without altering any logical structures, thereby enhancing the network’s generalization and robustness at a minimal additional cost. Extensive experiments indicate that HAIR possesses strong universality.
  • To the best of our knowledge, our method is the first to employ data-adaptive Hypernetworks in image restoration tasks, demonstrating their strong potential and offering a novel perspective for advancing image restoration capabilities.
@@ 2.1 All-in-One Image Restoration
While single degradation methods do achieve great success [9, 10, 11, 7, 12, 13], blind all-in-one image restoration, which aims to utilize a single deep blind restoration model to tackle multiple degradation types simultaneously, gaining more attention recently  [1, 2, 30, 24, 25, 26]. The seminal work AirNet [20] achieves blind All-in-One image restoration by employing contrastive learning to extract degradation representations from corrupted images, which are subsequently utilized to restore the clean image. Next, IDR [25] decomposes image degradations into their underlying physical principles, achieving All-in-One image restoration through a two-stage process based on meta-learning. With the rise of LLMs [31], Prompt-based learning [24, 27, 14] also emerged as a promising direction in computer vision tasks. Despite the success these methods achieved, they typically use the same parameters for distinct degradations and inject the degradation information as conditional embeddings into the model, which can impair the performance of the models. In contrast, our proposed HAIR can adaptively generate specific parameters for the given degraded image, resulting in better generalization to unknown degradations. [32] explains in theory why using Data-conditional Hypernetworks to generate adaptive parameters work better than injecting conditional embeddings into the model.
2.2 Image Classification
Image Classification is a fundamental task in Computer Vision, with the goal of categorizing images based on their content [33]. Typically, the process of Image Classification involves two key steps: feature extraction and the subsequent classification itself. The primary challenge in this task is feature extraction, which involves converting an image into a vector that encapsulates its most critical information. While traditional methods relied on handcrafted techniques [34], the advent of neural networks has revolutionized feature extraction [35, 36, 37, 38]. Given that these vectors are capable of classifying hundreds of categories [39], they must contain the essential information of the images. This concept extends to the classification of degraded images as well. For instance, if we categorize degraded images into types such as noisy, rainy, or hazy, and by levels of degradation like slightly noisy, moderately noisy, and very noisy, the same vectors should be effective in classifying these variations. The rationale is that recognizing different degradations is generally less complex than identifying hundreds of distinct objects. Therefore, our proposed method employs a Classifier to obtain a Global Information Vector (GIV). This GIV, rich in the global information of the image, including its degradation characteristics, is then utilized to generate parameters tailored to the specific degradation present in the image.
@@ 2.3 Data-conditioned Hypernetworks
Hypernetworks [29] are a class of neural networks designed to generate weight values (parameters) for other networks. They can be broadly classified into three main types: Task-conditioned, Data-conditioned, and Noise-conditioned hypernetworks [40]. Of these, Data-conditioned hypernetworks are particularly noteworthy for their ability to generate weights contingent upon the distinctive features of the input data. This capability allows the neural network to dynamically adapt its behavior in response to specific input patterns or characteristics, fostering flexibility and adaptability within the model. Consequently, this results in enhanced generalization and robustness. Data-conditioned Hypernetworks have found applications across the field of computer vision, including semantic segmentation [41] and image editing [42]. Despite previous attempts to integrate Hypernetworks into Image Restoration [43], these have primarily leveraged Task-conditioned hypernetworks. Such an approach can’t fully harness the potential of hypernetworks for the task of image restoration. To the best of our knowledge, this work pioneers the introduction of Data-conditioned hypernetworks to the domain of image restoration, unlocking new possibilities for the field.
Figure 3: The overall framework of HAIR. In this paper, we incorporate HAIR into Restormer [12], which has demonstrated robust performance in all-in-one restoration scenarios [24, 26, 28]. HAIR introduces a Classifier that yields a Global Information Vector (GIV) $\mathbf{V_{g}}$ from the high-level feature $\mathbf{F_{g}}$ encapsulating the input image’s degradation information. Additionally, the Hyper Selecting Nets (HSNs) dynamically generate weights for the Transformer Blocks based on GIV, customizing the restoration process to address the specific degradation characteristics of each image. This innovative approach allows models to adapt to various image restoration tasks without altering the original architecture.
@@ 3 Method
In this section, we outline the core principles of our proposed effective all-in-one image restoration method.
Overall Pipeline As shown in Fig. 3, our network architecture is consistent with Restormer [12]. Given a certain degraded image input $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, HAIR utilizes a $3\times 3$ convolution to transform $\mathbf{X}$ into feature embeddings $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$, where $C$ denotes the number of channels. These feature embeddings then proceed through a 4-level hierarchical encoder-decoder, resulting in deep features $\mathbf{F_{r}}$. Each level of the encoder-decoder incorporates several Transformer blocks, with an increasing number from the top to the bottom level, ensuring computational efficiency. The left three blue "Transformer Blocks" function as an Encoder, designed to extract features from $\mathbf{F_{0}}$ and ultimately produce a global feature map $\mathbf{F_{g}}\in\mathbb{R}^{\frac{W}{8}\times\frac{H}{8}\times 8C}$ with a large receptive field. $\mathbf{F_{g}}$ is then input into a Classifier to yield a Global Information Vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$. The right four orange "HyperTrans Blocks" operate as a Decoder, aiming to adaptively fuse features at each Decoder level based on degradation, culminating in a restored image $\hat{\mathbf{X}}$. Specifically, at each Decoder level, the HyperTrans Block accepts two inputs: the input feature and the GIV $\mathbf{V_{g}}$. $\mathbf{V_{g}}$ is fed into the Hyper Selecting Net to generate weights for the corresponding Transformer Blocks, which are then applied to the input feature to produce the output. It is important to note that all HyperTrans Blocks use the same $\mathbf{V_{g}}$ derived from $\mathbf{F_{g}}$. Since the weights in the Decoder are generated based on $\mathbf{V_{g}}$, which contains the degradation information of $\mathbf{X}$, our method exhibits strong generalization and adaptability, leading to excellent restoration performance.
@@ 3.1 Classifier
As outlined in Section 2.2, our approach to extracting degradation information involves designing a straightforward Classifier akin to those used in image classification tasks. We commence by inputting $\mathbf{F_{g}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ into the backbone, which is comprised of multiple ResNet Blocks [35] followed by ×$\times 2$ Downsampling. This process is aimed at downsizing the spatial resolution while distilling the essential information, as depicted by the following equations:
@@ 3.2 Hyper Selecting Net
$\mathbf{V_{s}}=\sigma(\mathbf{FCNN}(\mathbf{V_{g}}))$ (3)
Here, $\sigma$ stands for a Softmax operation and $\mathbf{FCNN}$ denotes a simple Fully-Connected Neural Network. Subsequently, the parameters $\mathbf{w}$ are derived:
$\mathbf{w}=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}$ (4)
@@
$\mathbf{x}^{\prime}=\mathbf{Transformer\_Block}(\mathbf{x};\mathbf{w})$ (5)
@@ 3.3 Rethinking Weights Generation via a Selecting Perspective
Proposition 1
Convolution operations exhibit the distributive law over addition. Let $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ be the input feature, and let $w_{i},\ i=1,2,\cdots,n$ represent the convolution kernels. The law is mathematically expressed as:
@@
With this proposition, the following becomes evident:
$\mathbf{x}*\mathbf{w}$ $=\mathbf{x}*\left(\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}\right)$ (7)
$=\sum_{i=1}^{N}\left(\mathbf{x}*(\mathbf{V_{s}}^{i}\mathbf{W}_{i})\right)$
$=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\left(\mathbf{x}*\mathbf{W}_{i}\right)$
With the constraint $\sum_{i}^{N}\mathbf{V_{s}}^{i}=1$ due to the property of Softmax operation. In essence, the convolution with the generated weights is equivalent to a weighted sum of convolutions between the input $\mathbf{x}$ and each kernel $\mathbf{W}_{i}$ in the Weights Box. This process can be seen as a process of selecting convolution kernels. For instance, with a Weights Box $\mathbf{W}\in\mathbb{R}^{2\times P}$ that includes a low-pass $\mathbf{W}_{1}$ and a high-pass $\mathbf{W}_{2}$, the rainy input $\mathbf{x}_{1}$ would ideally have a larger $\mathbf{V_{s}}^{1}$ to filter out high-frequency noise, while a smaller $\mathbf{V_{s}}^{2}$ would help retain details. Conversely, for a hazy input $\mathbf{x}_{2}$, a larger $\mathbf{V_{s}}^{2}$ would be necessary to mitigate low-frequency haze. Although real-world degradations can be more complex, our Hyper Selecting Net is adept at selecting the appropriate weights. This insight into the selection process helps us understand HAIR: for each input, the Hyper Selecting Net produces a tailored Selecting Vector and adaptively chooses the most suitable convolution kernel. Given that core operations such as convolution and matrix multiplication follow the distributive law over addition, HAIR is versatile and can be integrated into more various architectures like Transformer [46] and Mamba [47].

### v2

@@ Abstract
Image restoration aims to recover a high-quality clean image from its degraded version. Recent progress in image restoration has demonstrated the effectiveness of All-in-One image restoration models in addressing various degradations simultaneously. However, these existing methods typically utilize the same parameters to tackle images with different degradation types, thus forcing the model to balance the performance between different tasks and limiting its performance on each task. To alleviate this issue, we propose HAIR, a Hypernetworks-based All-in-One Image Restoration method that dynamically generates parameters based on input images. Specifically, HAIR consists of two main components, i.e., Classifier and Hyper Selecting Net (HSN). The Classifier is a simple image classification network used to generate a Global Information Vector (GIV) that contains the degradation information of the input image, and the HSN is a simple fully-connected neural network that receives the GIV and outputs parameters for the corresponding modules. Extensive experiments demonstrate that HAIR can significantly improve the performance of existing image restoration models in a plug-and-play manner, both in single-task and all-in-one settings. Notably, our innovative model, Res-HAIR, which integrates HAIR into the well-known Restormer, can obtain superior or comparable performance compared with current state-of-the-art methods. Moreover, we theoretically demonstrate that our proposed HAIR requires fewer parameters in contrast to the prevalent All-in-One methodologies. The code is available at https://github.com/toummHus/HAIR.
@@ 1 Introduction
Image restoration stands as a pivotal challenge within the field of computer vision, dedicated to the reconstruction of high-quality images from their degraded states. The presence of adverse conditions such as noise, haze, or rain can severely diminish the effectiveness of images for a variety of applications, such as autonomous navigation [1, 2], augmented reality [3, 4, 5]. Therefore, developing robust image restoration methods is of great importance. The use of deep learning in this domain has made remarkable progress, as evidenced by a suite of recent methodologies [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Nonetheless, the predominant approach in current research is to employ task-specific models, each tailored to address a particular type of degradation [16, 7, 12, 13, 8]. This tailored approach, while precise, presents a constraint in terms of universality, as it restricts the applicability of models to scenarios with varied or unknown degradations [17, 18, 19]. To overcome this limitation, the research community has increasingly turned its attention to the development of All-in-One image restoration models. These models are designed to tackle various degradations using one single model. Pioneering efforts in this area [20, 21, 1, 22, 23, 24, 25, 26, 15] have utilized a variety of advanced techniques such as contrastive learning [20], meta-learning [25], visual prompting methods [24, 27, 28, 14]. These approaches have undoubtedly made substantial contributions to the field, propelling it forward with notable advancements.
@@
Figure 1: The difference between our method and previous methods. (a) Previous All-in-One image restoration methods. For images with various degradations, these methods utilize a single model with the same parameters. (b) Our proposed HAIR. Given a certain degraded image, we first use a fixed Extraction module to extract the features, and feed the extracted feature into the Hypernetwork to generate the dynamic parameters for the fusion module and finally get the restored image by the generated parameters.
@@
Despite the notable successes of these methods, they share a common drawback, namely, they rely on a single model with fixed parameters to address a variety of degradations. This one-size-fits-all approach can hinder the model’s effectiveness when dealing with multiple degradations simultaneously. For example, when viewed through the lens of frequency domain analysis, haze is characterized as low-frequency noise, in contrast to rain, which is considered high-frequency interference. An effective dehazing model acts as a low-pass filter, preserving high-frequency details, whereas deraining requires the opposite—enhancing the high-frequency components. Consequently, a model must strike a balance between these conflicting demands of different degradations, which may lead to suboptimal performance.
@@
To mitigate the aforementioned issue, we propose HAIR in this paper. Our fundamental concept is the dynamic generation of parameters (weights) based on the degradation information present in the input image. HAIR employs Hypernetworks [29], a trainable neural network that outputs parameters for other neural networks, to take the degradation information from the input image and produce corresponding parameters. Specifically, for a given unknown degraded image, we first utilize a Classifier, similar to those used in image classification networks, to obtain its Global Information Vector, which contains crucial discriminative information about various types of image degradation (as shown in Fig. 2). This vector is then passed to the Hyper Selecting Net, which is a Fully Connected Network, to generate the requisite parameters, as illustrated in Fig. 1. Therefore, our method can dynamically adapt to various degradations. With these dynamically parameterised modules, we ultimately achieve the restored image. In brief, our contributions include:
  • We propose HAIR, a novel Hypernet-based All-in-One image restoration method capable of dynamically generating parameters based on the input image’s characteristics. HAIR consists of two components, i.e. Classifier and Hyper Selecting Nets, both of which function as a plug-in-and-play module. Extensive experiments demonstrate that HAIR can significantly improve the performance of existing image restoration models in a plug-and-play manner, both in single-task and all-in-one settings.
  • By incorporating HAIR into Restormer [12], we propose a new all-in-one model, i.e. Res-HAIR. To the best of our knowledge, our method is the first to apply data-adaptive Hypernetworks to All-in-One image restoration models. Extensive experiments validate that the proposed Res-HAIR can achieve superior or comparable performance compared with current state-of-the-art methods across a variety of image restoration tasks.
  • We theoretically prove that, for a given small error threshold $\epsilon$ in image restoration tasks, HAIR requires fewer parameters compared to mainstream embedding based All-in-One methods like [20, 24, 26].
@@ 2.1 All-in-One Image Restoration
While single degradation methods do achieve great success [9, 10, 11, 7, 12, 13], blind all-in-one image restoration, which aims to utilize a single deep blind restoration model to tackle multiple types of degradation simultaneously without the prior information of the degradation of the input image, has gained more attention recently  [1, 2, 30, 24, 25, 26]. The seminal work AirNet [20] achieves blind All-in-One image restoration by employing contrastive learning to extract degradation representations from corrupted images, which are subsequently utilized to restore the clean image. Next, IDR [25] decomposes image degradations into their underlying physical principles, achieving All-in-One image restoration through a two-stage process based on meta-learning. With the rise of LLM [31], prompt-based learning [24, 27, 14] has also emerged as a promising direction in computer vision tasks. Despite the success these methods have achieved, they typically use the same parameters for distinct degradations and inject the degradation information as conditional embeddings into the model, which may impair the performance of the models. In contrast, our proposed HAIR can adaptively generate specific parameters for the given degraded image, resulting in better generalization to unknown degradations. [32] theoretically explains why using data-conditional Hypernetworks to generate adaptive parameters works better than injecting conditional embeddings into the model.
2.2 Image Classification and Feature Extraction
Image classification is a fundamental task in Computer Vision, with the goal of categorizing images based on their content [33]. Typically, image classification contains two key steps: feature extraction and classification. The primary challenge in this task is feature extraction, which involves converting an image into a vector that encapsulates its most critical information. While traditional methods relied on handcrafted techniques [34], the advent of neural networks has revolutionized feature extraction [35, 36, 37, 38]. Given that these vectors are capable of classifying hundreds of categories [39], they must contain the essential information of the images. This concept also extends to the classification of degraded images. For instance, if we categorize degraded images into different types, e.g. noisy, rainy, or hazy, and by levels of degradation like slightly noisy, moderately noisy, and very noisy, the feature vectors should be effective in classifying these variations. The rationale is that recognizing different degradations is generally less complex than identifying hundreds of distinct objects. Therefore, our proposed method employs a Classifier to obtain a Global Information Vector (GIV), that captures the rich global information of the image, including its degradation characteristics. (As shown in Fig. 2) This GIV is then utilized to generate parameters tailored to the specific degradation present in the image.
@@ 2.3 Data-conditioned Hypernetworks
Hypernetworks [29] are a class of neural networks designed to generate weight values (parameters) for other networks. They can be broadly classified into three main types: Task-conditioned, Data-conditioned, and Noise-conditioned hypernetworks [40]. Among these methods, Data-conditioned hypernetworks are particularly noteworthy for their ability to generate weights contingent upon the distinctive features of the input data. This capability allows the neural network to dynamically adapt its behaviour in response to specific input patterns or characteristics, fostering flexibility and adaptability within the model. Consequently, this results in enhanced generalization and robustness. Data-conditioned Hypernetworks have found applications across the field of computer vision, including semantic segmentation [41] and image editing [42]. Despite previous attempts to integrate Hypernetworks into Image Restoration [43], these have primarily leveraged Task-conditioned hypernetworks. Such an approach cannot fully harness the potential of hypernetworks for the task of image restoration. To the best of our knowledge, our work first introduces Data-conditioned hypernetworks to the domain of image restoration, unlocking new possibilities for the field.
Figure 3: The overall framework of Res-HAIR. In this paper, we mainly incorporate HAIR into Restormer [12], which has demonstrated robust performance in all-in-one restoration scenarios [24, 26, 28]. HAIR introduces a Classifier that yields a Global Information Vector (GIV) $\mathbf{V_{g}}$ from the high-level feature $\mathbf{F_{g}}$ encapsulating the input image’s degradation information. Additionally, the Hyper Selecting Nets (HSNs) dynamically generate weights for the Transformer Blocks based on GIV, customizing the restoration process to address the specific degradation characteristics of each image. This innovative approach allows models to adapt to various image restoration tasks without altering the original architecture. For other baseline models, we follow a similar integration approach with HAIR: we introduce a classifier at the network’s midpoint and utilize the resulting GIV to dynamically generate weights for the subsequent layers of the network.
@@ 3 Method
In this section, we outline the core principles of our proposed effective all-in-one image restoration method. Note that we use Res-HAIR stands for Restormer embedded with HAIR.
Overall Pipeline As shown in Fig. 3, our network architecture is consistent with Restormer [12]. Given a certain degraded image input $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, Res-HAIR utilizes a $3\times 3$ convolution to transform $\mathbf{X}$ into feature embeddings $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$, where $C$ denotes the number of channels. These feature embeddings are then proceeded through a 4-level hierarchical encoder-decoder, resulting in deep features $\mathbf{F_{r}}$. Each level of the encoder-decoder incorporates several Transformer blocks, with an increasing number from the top to the bottom level, ensuring computational efficiency. The left three blue "Transformer Blocks" function as an Encoder, designed to extract features from $\mathbf{F_{0}}$ and ultimately produce a global feature map $\mathbf{F_{g}}\in\mathbb{R}^{\frac{W}{8}\times\frac{H}{8}\times 8C}$ with a large receptive field. $\mathbf{F_{g}}$ is then input into a Classifier to yield a Global Information Vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$. The right four orange "HyperTrans Blocks" operate as a Decoder, aiming to adaptively fuse features at each Decoder level based on degradation, culminating in a restored image $\hat{\mathbf{X}}$. Specifically, at each Decoder level, the HyperTrans Block accepts two inputs: the input feature and the GIV $\mathbf{V_{g}}$. $\mathbf{V_{g}}$ is fed into the Hyper Selecting Net to generate weights for the corresponding Transformer Blocks, which are then applied to the input feature to produce the output. It is important to note that all HyperTrans Blocks use the same $\mathbf{V_{g}}$ derived from $\mathbf{F_{g}}$. Since the weights in the Decoder are generated based on $\mathbf{V_{g}}$, which contains the degradation information of $\mathbf{X}$, our method exhibits strong generalization and adaptability, leading to excellent restoration performance.
@@ 3.1 Classifier
As outlined in Section 2.2, our approach to extracting degradation information involves designing a straightforward Classifier akin to those used in image classification tasks. We start by entering $\mathbf{F_{g}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ into the backbone, which is composed of multiple ResNet blocks [35] followed by downsampling ×$\times 2$. This process is aimed at downsizing the spatial resolution while distilling the essential information, as depicted by the following equations:
@@ 3.2 Hyper Selecting Net
$\mathbf{V_{s}}=\sigma(\mathbf{FCNN}(\mathbf{V_{g}})).$ (3)
Here, $\sigma$ stands for a Softmax operation and $\mathbf{FCNN}$ denotes a simple Fully-Connected Neural Network. Subsequently, the parameters $\mathbf{w}$ are derived as
$\mathbf{w}=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}.$ (4)
@@
$\mathbf{x}^{\prime}=\mathbf{Transformer\_Block}(\mathbf{x};\mathbf{w}).$ (5)
@@ 3.3 Rethinking Weights Generation via a Selecting Perspective
Proposition 1.
(Convolution operations exhibit the distributive law over addition) Let $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ be the input feature, and let $w_{i},\ i=1,2,\cdots,n$ represent the convolution kernels. The law is mathematically expressed as:
@@
With this proposition, the following formula becomes evident:
$\mathbf{x}*\mathbf{w}=\mathbf{x}*\left(\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}\right)=\sum_{i=1}^{N}\left(\mathbf{x}*(\mathbf{V_{s}}^{i}\mathbf{W}_{i})\right)=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\left(\mathbf{x}*\mathbf{W}_{i}\right),$ (7)
where the constraint $\sum_{i}^{N}\mathbf{V_{s}}^{i}=1$ holds due to the property of Softmax operation. In essence, the convolution with the generated weights is equivalent to a weighted sum of convolutions between the input $\mathbf{x}$ and each kernel $\mathbf{W}_{i}$ in the Weights Box. This process can be seen as a process of selecting convolution kernels. For instance, with a Weights Box $\mathbf{W}\in\mathbb{R}^{2\times P}$ that includes a low-pass $\mathbf{W}_{1}$ and a high-pass $\mathbf{W}_{2}$, the rainy input $\mathbf{x}_{1}$ would ideally have a larger $\mathbf{V_{s}}^{1}$ to filter out high-frequency noise, while a smaller $\mathbf{V_{s}}^{2}$ would help retain details. Conversely, for a hazy input $\mathbf{x}_{2}$, a larger $\mathbf{V_{s}}^{2}$ would be necessary to mitigate low-frequency haze. Although real-world degradations can be more complex, our Hyper Selecting Net can adaptively select the appropriate weights. This insight into the selection process helps us understand HAIR, namely, for each input, the Hyper Selecting Net produces a tailored Selecting Vector and adaptively chooses the most suitable convolution kernel. Given that core operations such as convolution and matrix multiplication follow the distributive law over addition, HAIR is versatile and can be integrated into various architectures like Transformer [46] and Mamba [47].